# A revisited Correction to the Halo Mass Function for local-type Primordial non-Gaussianity

### v1

@@ 2.2 Non-Gaussian models for the halo mass function
Finally, we introduce the key approximation that these correction factors are independent of the specific model used to describe the mass function. Consequently, once the correction factor $\mathcal{R}(M,z)$ is determined, it can be applied to any Gaussian model $n(M,z)$, such as those proposed in [51, 52, 53, 54], to obtain the desired non-Gaussian model:
@@ 3.2 Data-theory comparison
Figure 1: The average Gaussian halo mass function measured from 8000 realizations of the quijote simulations at redshift $z=0$, within the mass range of $5\times 10^{13}$ to $7\times 10^{15}$ $h^{-1}\ \mathrm{M}_{\odot}$. The purple, green and orange circles correspond to different mass definitions: $M_{\mathrm{200b}}$, $M_{\mathrm{200c}}$, and $M_{\mathrm{500c}}$, respectively. The solid lines indicate the theoretical model proposed by Tinker et al. [67] for each definition of mass. The subpanel displays the percentage residuals of the data relative to the model.
We also compare the estimated mass functions with the theoretical model proposed by Tinker et al. [67], which can be expressed as:
@@
As illustrated in the subplot of Figure 1, the theoretical model does not reproduce the data with high accuracy. For halos defined by $M_{\mathrm{200b}}$, the agreement is relatively good, with residuals around $5\%$ (excluding the first two data points). However, as we move to other mass definitions, the agreement deteriorates, with residuals increasing up to 10% and 20% (in absolute value) for $M_{\mathrm{200c}}$ and $M_{\mathrm{500c}}$, respectively. Similar results are also observed when the mass functions are examined at higher redshifts. We repeated the analysis on the halo mass functions extracted from pringls and verified its consistency with the quijote data in all definitions of mass and redshifts. Also in this case the theoretical model again showed poor agreement with the pringls data, despite the larger error bars due to the smaller number of available realizations (only $10$ per cosmological model).
@@
Figure 2: Deviations between halo mass functions in PNG scenarios with respect to their Gaussian counterpart. We compare different theoretical models and data sets at $z=0$, for the two scenarios $f_{\mathrm{NL}}=100$ and $-100$. On the $y$-axis we show the percentage residuals (Eq. 3.2) between the mean halo number counts measured using quijote-png and those extracted from quijote (representing the reference $\Lambda\mathrm{CDM}$ number counts). Different marker shapes are used for the two $f_{\mathrm{NL}}$ values: upward-pointing triangles for $+100$ and downward-pointing triangles for $-100$. Colors denote different mass definitions: green for $M_{\mathrm{200b}}$, blue for $M_{\mathrm{200c}}$, and orange for $M_{\mathrm{500c}}$. The theoretical models for mass function corrections are indicated with different line styles: solid for MVJ, dashed for LMSVQ, and dotted for DMNP (Eqs. 2.10, 2.12 and 2.13, respectively). The subplot displays a zoomed region at low masses to allow a clearer understanding of the behavior of both model and data points.
@@ 3.3 Density profiles of dark matter halos
DM halos exhibit an increase in inner density for positive values of $f_{\mathrm{NL}}$, while a decrease is observed when $f_{\mathrm{NL}}$ is negative. By construction, all profiles become identical when the x-axis reaches unity; in fact, because of the mass definition used, every halo is characterized by the same density contrast at a distance $R=R_{\mathrm{200b}}$ from the center. Furthermore, the measured mean deviations from the Gaussian profile appear symmetric, with variations of at most 2% in the innermost regions of the halos. Within each mass bin, the number of halos increases as $f_{\mathrm{NL}}$ increases, with the opposite behavior occurring as $f_{\mathrm{NL}}$ decreases, as shown in Table 2.
The change we observe in the slope of the density profiles, i.e. in the compactness of DM halos, is reflected in the deviations found in the halo mass function for different values of the identification threshold (see Figure 2). In fact, given a threshold $\Delta$, the radius at which that density contrast is reached depends on the value of $f_{\mathrm{NL}}$, and the mass of the halo changes as a consequence. For example, for positive values of $f_{\mathrm{NL}}$, the threshold is met at a greater distance from the halo center, resulting in a larger mass within the corresponding sphere. Additionally, higher $\Delta$ values cause the identified halo radii to be closer to the center, leading to greater deviations from the Gaussian case. Our goal is to incorporate these effects into the theoretical model of the halo mass function by adding a corrective factor to account for the variations caused by the PNG in halo density profiles.
@@ 4.1 Calibration of the correction factor $\kappa$
Based on the results presented in the previous sections, it is clear that the theoretical models, without adjustments for specific mass definitions, do not adequately agree with data from simulations, especially at higher overdensities. To address this, we sought a method that incorporates the effects of varying mass definitions. In the literature, it is common to introduce a correction factor to align theoretical models with FoF data sets when correcting the halo mass function for PNG (as discussed in Section 3.2). We adopted a similar approach by introducing a modification of the linear critical density for collapse, parameterized by a factor $\kappa$, which varies with the chosen overdensity criterion used to identify the halos. Among all the theoretical non-Gaussian mass function models analyzed, we chose to rely on the LMSVQ model since it showed the best agreement with our data (see e.g. the $M_{\mathrm{200b}}$ case in Figure 2). Consequently, we modify Eq. 2.12 by substituting $\delta_{\mathrm{c}}$ with $\kappa\times\delta_{\mathrm{c}}$.
@@
Figure 4: Left: factor $\kappa$ as a function of the overdensity $\Delta_{\mathrm{b}}$. Different markers indicate the two opposite values of the $f_{\mathrm{NL}}$ parameter in the quijote-png simulations (upward-pointing triangles for $+100$ and downward-pointing triangles for $-100$), while colors correspond to redshifts $z=0$, $0.5$, and $1$ (cyan, light purple, and magenta, respectively). In red and blue we indicate the $1\sigma$ confidence region around the best fit of the $f_{\mathrm{NL}}=100$ and $-100$ data points, respectively (note that these regions are almost perfectly overlapping). The dark purple curve represents the simultaneous fit of both data sets, including the effect of covariance. Right: 68% and 95% 2D confidence contours for the parameters $a$, $b$, and $c$ characterizing the second-degree polynomials (Eq. 4.2) for each fitting procedure. The same colors are used to indicate the results for $f_{\mathrm{NL}}=100$, $-100$ and their combination. The projected 1D marginalized posterior distributions are shown at the top of each column along with the 68% uncertainty (shaded bars).
@@
Figure 5 presents the covariance matrix C normalized by the diagonal elements, estimated from the residuals between the individual $2\times 500$ halo mass functions $M_{\mathrm{200b}}$ measured in each realization of quijote-png ($500$ for $f_{\mathrm{NL}}=100$ and $500$ for $f_{\mathrm{NL}}=-100$) and the mean halo mass function extracted from quijote ($f_{\mathrm{NL}}=0$). Similar covariance matrices are also obtained for the other mass definitions. Each element of these matrices is rescaled by a factor of $500$ to account for the number of realizations; in this way, the diagonal elements will correctly represent the error on the mean of the residuals. Figure 5 also shows a strong correlation between the same mass bins in the $f_{\mathrm{NL}}=\pm 100$ data sets, whereas for the other bins, the correlation is consistent with zero, except for minor fluctuations due to noise. To avoid possible inaccuracies due to the latter, we set to zero all elements that do not belong to the diagonals of the four blocks in each matrix.
@@ 4.2 Testing different $f_{\mathrm{NL}}$ values
Figure 7: $68\%$ and $95\%$ confidence levels for the coefficient $a$, $b$, and $c$ of the second-degree polynomial used to fit the values of $\kappa(\Delta_{\mathrm{b}})$ derived with pringls simulations. Light blue, blue, red and orange correspond to the different values of $f_{\mathrm{NL}}$: $-100$, $-40$, $40$, $100$, respectively. Black dashed lines are the best fit values of the parameters reported in Eq. 4.4, which we obtained using quijote-png. The projected 1D marginalized posterior distributions are shown at the top of each column along with the 68% uncertainty (shaded bars).
@@
Most importantly, given the compatibility of the results obtained for all the scenarios analyzed, we can conclude that the calibrated relation for $\kappa(\Delta_{\mathrm{b}})$ is not only independent of the sign of $f_{\mathrm{NL}}$, but also of its absolute value. This outcome is clearly very important: given a sample of DM halos identified with any threshold $\Delta$, we now have a unique and simple recipe to predict the mass function in scenarios featuring whatever level of local-type PNG.
As a final consideration, we note that the proposed re-parametrization of $\delta_{\mathrm{c}}$ by means of the factor $\kappa(\Delta_{\mathrm{b}})$ can be applied also to other non-Gaussian halo mass function models, for example MVJ [36] and DMNP [39]. We confirm that this modification is indeed effective in correcting the model predictions according to the selected halo density thresholds, although the best agreement with the data is achieved using the LSMVQ model. Since we used the latter to calibrate the $\kappa(\Delta_{\mathrm{b}})$ relation, we recommend applying Eq. 4.4 solely in combination with the LMSVQ model for an accurate prediction of the non-Gaussian halo mass function.
@@ 5 Discussions and Conclusions
In this paper, we focused on testing the accuracy of already existing halo mass function models developed for cosmological scenarios featuring primordial non-Gaussianities (PNG). We selected three notorious models from the literature, that is, Matarrese et al. (2000) [36], LoVerde et al. (2008) [38], and D’Amico et al. (2011) [39], to test against simulations. These models provide a theoretical prescription to predict the deviation from the Gaussian halo mass function produced by local type of PNG, quantified by the parameter $f_{\mathrm{NL}}$. The data were extracted from halo catalogs generated by applying the algorithm rockstar [62] to the three sets of N-body cosmological simulations quijote, quijote-png and pringls (see Section 3.1). These simulations share the same standard $\Lambda$CDM cosmological parameters and together cover different levels of PNG, i.e. $f_{\mathrm{NL}}=0$, $f_{\mathrm{NL}}=\pm 40$ and $f_{\mathrm{NL}}=\pm 100$. A key point in our analysis consisted in the creation of catalogs with different thresholds $\Delta$ for halo identification. In particular, we focused on different threshold values, defined both with respect to $\rho_{\mathrm{b}}$ (e.g. $\Delta_{\mathrm{b}}=200$) and $\rho_{\mathrm{c}}$ (e.g. $\Delta_{\mathrm{c}}=200$, $500$, $2500$).
@@
Finally, we underline that the methodology proposed in this paper is not meant to be limited only to local-type PNG: a natural extension of this work will consist of testing the models present in the literature also for other shapes of the potential bispectrum (i.e. equilateral, enfolded, orthogonal). Moreover, one of our goals is to extend our analysis to the underdense counterpart of DM halos, namely cosmic voids. In fact, the same formalism used to predict the non-Gaussian halo mass function can be applied to void counts, allowing us to model the non-Gaussian void size function. These two models can ultimately be used in combination to extract the overall information from halo and void number counts. In fact, several studies have demonstrated the strong complementarity of these two statistics [see e.g. 74, 75, 76, 77], making them excellent probes for breaking key degeneracies between cosmological parameters.
@@ Appendix A Non-Gaussian Halo Bias
where $\delta_{\mathrm{c,0}}=1.686$ and $p$ is a parameter introduced to take into account possible deviations caused by the non-universality of the halo mass function model. In particular, the assumption of a universal mass function leads to $p=1$. This expression for $p$ is often referred to as the universal relation for $b_{\Phi}$.
However, recent studies suggest that this relation does not always accurately describe the scale-dependent bias. For example, [84] found that for galaxies selected by stellar mass, $p=0.5$ provides a better fit for the $b_{\Phi}$ relation. Similarly, in [85], an analysis of halos within a specific mass range found $p=0.955\pm 0.013$, indicating a discrepancy of about $3\sigma$ from the universal relation for $b_{\Phi}$. Moreover, $b_{\Phi}$ may depend on additional properties beyond mass, such as halo concentration [86, 87].

### v2

@@ 2.2 Non-Gaussian models for the halo mass function
Finally, we introduce the key approximation that these correction factors are independent of the specific model used to describe the mass function. Consequently, once the correction factor $\mathcal{R}(M,z)$ is determined, it can be applied to any Gaussian model $n(M,z)$, such as those proposed in [51, 52, 53, 54, 55], to obtain the desired non-Gaussian model:
@@ 3.2 Data-theory comparison
Figure 1: The average Gaussian halo mass function measured from 8000 realizations of the Quijote simulations at redshift $z=0$, within the mass range of $5\times 10^{13}$ to $7\times 10^{15}$ $h^{-1}\ \mathrm{M}_{\odot}$. The purple, green and orange circles correspond to different mass definitions: $M_{\mathrm{200b}}$, $M_{\mathrm{200c}}$, and $M_{\mathrm{500c}}$, respectively. The solid lines indicate the phenomenological model proposed by Tinker et al. [55] for each definition of mass. The subpanel displays the percentage residuals of the data relative to the model.
We also compare the estimated mass functions with the phenomenological model proposed by Tinker et al. [55], which can be expressed as:
@@
As illustrated in the subplot of Figure 1, the model does not reproduce the data with high accuracy. For halos defined by $M_{\mathrm{200b}}$, the agreement is relatively good, with residuals around $5\%$ (excluding the first two data points). However, as we move to other mass definitions, the agreement deteriorates, with residuals increasing up to 10% and 20% (in absolute value) for $M_{\mathrm{200c}}$ and $M_{\mathrm{500c}}$, respectively. Similar results are also observed when the mass functions are examined at higher redshifts. We repeated the analysis on the halo mass functions extracted from Pringls and verified its consistency with the Quijote data in all definitions of mass and redshifts. Also in this case the phenomenological model again showed poor agreement with the Pringls data, despite the larger error bars due to the smaller number of available realizations (only $10$ per cosmological model).
@@
Figure 2: Deviations between halo mass functions in PNG scenarios with respect to their Gaussian counterpart. We compare different theoretical models and data sets at $z=0$, for the two scenarios $f_{\mathrm{NL}}=100$ and $-100$. On the $y$-axis we show the percentage residuals (Eq. 3.2) between the mean halo number counts measured using Quijote-png and those extracted from Quijote (representing the reference $\Lambda\mathrm{CDM}$ number counts). Different marker shapes are used for the two $f_{\mathrm{NL}}$ values: upward-pointing triangles for $100$ and downward-pointing triangles for $-100$. Colors denote different mass definitions: green for $M_{\mathrm{200b}}$, blue for $M_{\mathrm{200c}}$, and orange for $M_{\mathrm{500c}}$. The theoretical models for mass function corrections are indicated with different line styles: solid for MVJ, dashed for LMSVQ, and dotted for DMNP (Eqs. 2.10, 2.12 and 2.13, respectively). The subplot displays a zoomed region at low masses to allow a clearer understanding of the behavior of both model and data points.
@@ 3.3 Density profiles of dark matter halos
DM halos exhibit an increase in inner density for positive values of $f_{\mathrm{NL}}$, while a decrease is observed when $f_{\mathrm{NL}}$ is negative. By construction, all profiles become identical when the x-axis reaches unity; in fact, because of the mass definition used, every halo is characterized by the same density contrast at a distance $R=R_{\mathrm{200b}}$ from the center. Furthermore, the measured mean deviations from the Gaussian profile appear symmetric, with variations of at most 2% in the innermost regions of the halos. Within each mass bin, the number of halos increases as $f_{\mathrm{NL}}$ increases, with the opposite behavior occurring as $f_{\mathrm{NL}}$ decreases, as shown in Table 2. Interestingly, the bottom right panel of Figure 3 reveals that the $f_{\mathrm{NL}}=-40$ and $-100$ scenarios deviate less from the Gaussian case compared to their positive $f_{\mathrm{NL}}$ counterparts. This asymmetry suggests that the impact of negative amplitudes of PNG on the internal structure of high-mass halos is milder. However, this bin is characterized by the largest statistical uncertainty; therefore, we plan to investigate this aspect in future analyses using high-resolution simulations.
The change we observe in the slope of the density profiles, i.e. in the compactness of DM halos, is reflected in the deviations found in the halo mass function for different values of the identification threshold (see Figure 2). In fact, given a threshold $\Delta$, the radius at which that density contrast is reached depends on the value of $f_{\mathrm{NL}}$, and the mass of the halo changes as a consequence. For example, for positive values of $f_{\mathrm{NL}}$, the threshold is met at a greater distance from the halo center, resulting in a larger mass within the corresponding sphere. Additionally, higher $\Delta$ values cause the identified halo radii to be closer to the center, leading to greater deviations from the Gaussian case. Notably, the asymmetry observed at large masses in the density profiles is mirrored by a corresponding asymmetry in the deviations from the Gaussian case of the HMF, further supporting the interpretation that the modifications in halo structure due to PNG directly impact the abundance of halos. Our goal is to incorporate these effects into the theoretical model of the halo mass function by adding a corrective factor to account for the variations caused by the PNG in halo density profiles.
@@ 4.1 Calibration of the correction factor $\kappa$
Based on the results presented in the previous sections, it is clear that the theoretical models, without adjustments for specific mass definitions, do not adequately agree with data from simulations, especially at higher overdensities. To address this, we sought a method that incorporates the effects of varying mass definitions. In the literature, it is common to introduce a correction factor to align theoretical models with FOF data sets when correcting the halo mass function for PNG (as discussed in Section 3.2). We adopted a similar approach by introducing a modification of the linear critical density for collapse, parameterized by a factor $\kappa$, which varies with the chosen overdensity criterion used to identify the halos. Among all the theoretical non-Gaussian mass function models analyzed, we chose to rely on the LMSVQ model since it showed the best agreement with our data (see e.g. the $M_{\mathrm{200b}}$ case in Figure 2). Consequently, we modify Eq. 2.12 by substituting $\delta_{\mathrm{c}}$ with $\kappa\times\delta_{\mathrm{c}}$. Therefore, thanks to this additional factor, our model for the non-Gaussian HMF more closely resembles the structure of commonly used models for the Gaussian HMF, where the quantity $\delta_{\mathrm{c}}$ is typically multiplied by or replaced with a fitted parameter [see, e.g., 51, 52, 53, 54, 55].
@@
Figure 4: Left: factor $\kappa$ as a function of the overdensity $\Delta_{\mathrm{b}}$. Different markers indicate the two opposite values of the $f_{\mathrm{NL}}$ parameter in the Quijote-png simulations (upward-pointing triangles for $100$ and downward-pointing triangles for $-100$), while colors correspond to redshifts $z=0$, $0.5$, and $1$ (cyan, light purple, and magenta, respectively). In red and blue we indicate the $1\sigma$ confidence region around the best fit of the $f_{\mathrm{NL}}=100$ and $-100$ data points, respectively (note that these regions are almost perfectly overlapping). The dark purple curve represents the simultaneous fit of both data sets, including the effect of covariance. Right: 68% and 95% 2D confidence contours for the parameters $a$, $b$, and $c$ characterizing the second-degree polynomials (Eq. 4.2) for each fitting procedure. The same colors are used to indicate the results for $f_{\mathrm{NL}}=100$, $-100$ and their combination. The projected 1D marginalized posterior distributions are shown at the top of each column along with the 68% uncertainty (shaded bars).
@@
Figure 5 presents the covariance matrix C normalized by the diagonal elements, estimated from the residuals between the individual $2\times 500$ halo mass functions $M_{\mathrm{200b}}$ measured in each realization of Quijote-png ($500$ for $f_{\mathrm{NL}}=100$ and $500$ for $f_{\mathrm{NL}}=-100$) and the mean halo mass function extracted from Quijote ($f_{\mathrm{NL}}=0$). Similar covariance matrices are also obtained for the other mass definitions. Each element of these matrices is rescaled by a factor of $500$ to account for the number of realizations; in this way, the diagonal elements will correctly represent the error on the mean of the residuals. Figure 5 also shows a strong correlation between the same mass bins in the $f_{\mathrm{NL}}=\pm 100$ data sets, whereas for other bins, the correlation is consistent with zero, except for minor fluctuations due to noise. To avoid possible inaccuracies due to the latter, we set to zero all elements whose correlation coefficient lies between $-0.15$ and $0.15$.
@@ 4.2 Testing different $f_{\mathrm{NL}}$ values
Figure 7: $68\%$ and $95\%$ confidence levels for the coefficient $a$, $b$, and $c$ of the second-degree polynomial used to fit the values of $\kappa(\Delta_{\mathrm{b}})$ derived with Pringls simulations. Blue, light blue, orange and red correspond to the different values of $f_{\mathrm{NL}}$: $-100$, $-40$, $40$, $100$, respectively. Black dashed lines are the best fit values of the parameters reported in Eq. 4.4, which we obtained using Quijote-png. The projected 1D marginalized posterior distributions are shown at the top of each column along with the 68% uncertainty (shaded bars).
@@
Most importantly, given the compatibility of the results obtained for all the scenarios analyzed, we can conclude that the calibrated relation for $\kappa(\Delta_{\mathrm{b}})$ is not only independent of the sign of $f_{\mathrm{NL}}$, but also of its absolute value. We attribute this result to the fact that, in the expression for the halo mass function correction, all the information on the magnitude of $f_{\mathrm{NL}}$ is already well encapsulated in $S_{3}$ and $S_{4}$. Having no statistical evidence to the contrary, we conclude that it is not necessary to recalibrate the parameters of the Gaussian mass function according to the value of $f_{\mathrm{NL}}$. This outcome is clearly very important: given a sample of DM halos identified with any threshold $\Delta$, we now have a unique and simple recipe to predict the mass function in scenarios featuring whatever level of local-type PNG.
Furthermore, we note that the proposed re-parametrization of $\delta_{\mathrm{c}}$ by means of the factor $\kappa(\Delta_{\mathrm{b}})$ can be applied also to other non-Gaussian halo mass function models, for example MVJ [36] and DMNP [39]. We confirm that this modification is indeed effective in correcting the model predictions according to the selected halo density thresholds, although the best agreement with the data is achieved using the LSMVQ model. Since we used the latter to calibrate the $\kappa(\Delta_{\mathrm{b}})$ relation, we recommend applying Eq. 4.4 solely in combination with the LMSVQ model for an accurate prediction of the non-Gaussian halo mass function.
As a final consideration, we emphasize that, given the resolution of the simulations, it is not possible to totally exclude potential resolution-related effects, particularly on the estimate of the $\kappa$ parameter for halos identified with particularly high thresholds (such as $\Delta_{\mathrm{c}}=1200$ or greater). We refer the reader to Appendix B for a more in-depth analysis of this subject.
@@ 4.3 Impact on forecasted constraints
However, we stress that these constraints on $f_{\mathrm{NL}}$ are optimistic, as we have fixed all other cosmological parameters. There could be some degeneracies between the parameters characterizing our new version of the model and the cosmological ones. We aim to investigate this aspect in future analyses.
@@ 5 Discussions and Conclusions
In this paper, we focused on testing the accuracy of already existing halo mass function models developed for cosmological scenarios featuring primordial non-Gaussianities (PNG). We selected three well-known models from the literature, that is, Matarrese et al. (2000) [36], LoVerde et al. (2008) [38], and D’Amico et al. (2011) [39], to test against simulations. These models provide a theoretical prescription to predict the deviation from the Gaussian halo mass function produced by local type of PNG, quantified by the parameter $f_{\mathrm{NL}}$. The data were extracted from halo catalogs generated by applying the algorithm ROCKSTAR [63] to the three sets of N-body cosmological simulations Quijote, Quijote-png and Pringls (see Section 3.1). These simulations share the same standard $\Lambda$CDM cosmological parameters and together cover different levels of PNG, i.e. $f_{\mathrm{NL}}=0$, $f_{\mathrm{NL}}=\pm 40$ and $f_{\mathrm{NL}}=\pm 100$. A key point in our analysis consisted in the creation of catalogs with different thresholds $\Delta$ for halo identification. In particular, we focused on different threshold values, defined both with respect to $\rho_{\mathrm{b}}$ (e.g. $\Delta_{\mathrm{b}}=200$) and $\rho_{\mathrm{c}}$ (e.g. $\Delta_{\mathrm{c}}=200$, $500$, $2500$).
@@
Regarding this last point, although it is true that our results were obtained for relatively large values of $f_{\mathrm{NL}}$, we believe that similar improvements in terms of modeling systematics on the parameter would also be present when aiming to constrain more realistic values of $f_{\mathrm{NL}}$. However, in such cases the observable deviations would be significantly smaller than those considered here, and one would also need to account for the uncertainties in the number counts as well as those introduced, for instance, by the choice of the Gaussian mass function model. For this reason, we plan to investigate this aspect in future work.
Finally, we underline that the methodology proposed in this paper is not meant to be limited only to local-type PNG: a natural extension of this work will consist of testing the models present in the literature also for other shapes of the potential bispectrum (i.e. equilateral, enfolded, orthogonal). Moreover, one of our goals is to extend our analysis to the underdense counterpart of DM halos, namely cosmic voids. In fact, the same formalism used to predict the non-Gaussian halo mass function can be applied to void counts, allowing us to model the non-Gaussian void size function. These two models can ultimately be used in combination to extract the overall information from halo and void number counts. In fact, several studies have demonstrated the strong complementarity of these two statistics [see e.g. 77, 78, 79, 80], making them excellent probes for breaking key degeneracies between cosmological parameters.
@@ Appendix A Non-Gaussian Halo Bias
where $\delta_{\mathrm{c,0}}=1.686$ and $p$ is a parameter introduced to take into account possible deviations caused by the non-universality of the halo mass function model. In particular, the assumption of a universal mass function leads to $p=1$. This expression for $p$ is often referred to as the universal relation for $b_{\Phi}$. However, recent studies suggest that this relation does not always accurately describe the scale-dependent bias. For example, [87] found that for galaxies selected by stellar mass, $p=0.5$ provides a better fit for the $b_{\Phi}$ relation. Similarly, in [88], an analysis of halos within a specific mass range found $p=0.955\pm 0.013$, indicating a discrepancy of about $3\sigma$ from the universal relation for $b_{\Phi}$. Moreover, $b_{\Phi}$ may depend on additional properties beyond mass, such as halo concentration [89, 90].
@@
Appendix B Calibration against High Resolution Simulations
Defining halo masses within a radius enclosing a given overdensity $\Delta$ implies that the halo mass can be derived from the integral of the density profile within a fixed radius. Consequently, the mass depends on the internal density distribution of the halo and is therefore more sensitive to resolution effects, especially at high overdensities.
[…]
The results of this analysis are shown in Figure 9. We notice how the contours derived from the residuals for $f_{\mathrm{NL}}=100$ and $-100$ are fully compatible, highlighting that the independence of this new parameterization of the sign of $f_{\mathrm{NL}}$ is robust even when considering high-resolution simulations. Moreover, we observe from the posterior distributions that the results obtained with the Quijote-png data sets (vertical and horizontal dashed lines) are compatible within $1\sigma$ with those of Pringls-hr. Therefore, we conclude that our new parameterization does not seem to be significantly affected by resolution effects. However, further studies with additional realizations would help to confirm this prediction.
Figure 9: $68\%$ and $95\%$ confidence levels for the coefficients $a$, $b$, and $c$ of the second-degree polynomial used to fit $\kappa(\Delta_{\mathrm{b}})$ from Pringls-hr simulations. Light blue and orange contours correspond to $f_{\mathrm{NL}}=-100$ and $f_{\mathrm{NL}}=100$, respectively. Black dashed lines indicate the best-fit values of the parameters reported in Eq. 4.4, obtained using Quijote-png. The projected 1D marginalized posterior distributions are shown at the top of each column, with shaded bars representing the 68% uncertainty.